OpenStudy (anonymous):

How do I evaluate in the real number system: (0.027)^-2/3 / (0.25)^-1/2

OpenStudy (anonymous):

\[(0.027)^{-2/3}/(0.25)^{-1/2}\] = \[(^{-3}\sqrt{0.027})^{2} / (^{-2}\sqrt{0.25}) \] \[= (^{2}\sqrt{0.25}) / (^{3}\sqrt{0.027})^{2}\]

OpenStudy (anonymous):

idk if that's how I should do it and I'm stuck..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):